DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, March 13, 2002 Padelford C-36 3:45 pm

Propagation of Polarization in Elastodynamics

Sönke Hansen (Universität Paderborn)

Polarized light is a solution of Maxwell’s equations with a particular direction of oscillation. Maxwell’s equations for isotropic media are a system of real principal type. For such systems the polarization of solutions is governed by what is called the transport equation in geometrical optics. The Lamé equations of isotropic elastodynamics constitute another important example of a system of real principal type. Propagation of polarization is used for studying the inverse problem which aims at recovering elastic parameters. In the talk I report about the thesis of Daniel Röhrig and about joint work with Gunther Uhlmann.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.